### v2

@@ 3.3 Efficient GBP
Efficient inference is necessary for our models to be useful in practice. However, the inversion of a $\left(V_{j}-1\right)\times\left(V_{j}-1\right)$ matrix to compute $\Sigma_{\setminus i,\setminus i}^{\left(\phi_{j}+m\right)}$ in (8) has $O\left((V_{j}-1)^{3}\right)$ complexity, bottlenecking GBP. To alleviate this, we exploit the structure of the matrix being inverted. In particular, i) for factors with observation dimension $M=\dim\left(\mathbf{y}\right)<V_{j}$, the precision $\Lambda^{(\phi_{j})}$ (9) is low-rank, and ii) for graphs with scalar variable nodes, $\mathbf{D}$ is diagonal. Thus their sum may be efficiently inverted via the Woodbury identity (woodbury1950inverting). Further savings come from reusing intermediates when computing messages to multiple variables (9). These optimisations change the complexity of updating messages from a factor to all $V$ variables, from $O(V\left(V-1\right)^{3})$ to $O\left(VM^{3}\right)$. Space complexity is changed from $O(V^{2})$ to $O(VM+M^{2})$. See App. B for details. These results constitute significant savings when $M<<V$ which is usually true in practice.
@@ 4 Related Work
Our models can be viewed as probabilistic energy-based models (EBMs lecun2006tutorial; du2019implicit) whose energy functions are the sum of the quadratic energies for all factors in the graph. The benefit of using a Gaussian factor graph is a model which normalises in closed form without resorting to expensive MCMC sampling. While the quadratic energy may seem constraining, we add capacity with the introduction of non-linear factors and overparameterisation. For most EBMs, parameters are attributes of the factors, where we include them as variables in our graph. As a result, learning and prediction are the same procedure in our approach, but two distinct stages for other EBMs.
Of the EBM family, restricted Boltzmann machines (RBMs smolensky1986information) are of particular relevance. Their factor graph resembles a single layer, fully connected version of our model, however RBMs are models over discrete variables which are unable capture the statistics of continuous natural images. While exponential family generalisations exist, such as Gaussian-Bernoulli RBMs (welling2004exponential), scaling RBMs to multiple layers remains a challenging problem. To our knowledge there are no working examples of training stacked RBMs jointly, instead they are trained greedily, with each layer learning to reconstruct the activations of the trained and fixed previous layer (hinton2006fast; hinton2006reducing). As an artefact of this, later layers use capacity to learn artificial correlations induced by the early layers, rather than correlations present in the data. Global alignment is then found by fine tuning with either backprop (hinton2006reducing) or wake-sleep (hinton2006fast). We have found GBP Learning to work in multiple layer models without issue.
@@
Much work has focused on augmenting BP with DL (nachmani2016learning; satorras2021neural; yoon2019inference; lazaro2021query). In contrast, our method does DL within a BP framework. george2017generative propose a hybrid vision system in which visual features and graph structure are learnt in a separate process to the discrete BP routine used to parse scenes (given features and graph). lazaro2016hierarchical is more similar to ours in that parameters are treated as variables in the graphical model and updated with BP. However, their factor graph comprises only binary variables, making it ill-suited to natural images. In addition, max-product BP is used to find a MAP solution, and not a marginal posterior estimate. This precludes the ability to learn incrementally via Bayesian filtering.
@@ 5 Results
We demonstrate our approach with three experiments: some small regression tasks, sequential video denoising and image classification. Our TensorFlow (Tensorflow:MISC2015) implementation is made available222anonymous.4open.science/r/gbp-learning-anon-F2CB/.
@@
Figure 2: GBP Learning in MLP-like factor graphs (1(a)) can solve nonlinear regression and classification tasks. (1(b)) was generated with $8$ hidden units, (1(c)) with $16$ hidden units.
@@ 5.2 Video Denoising
To evaluate the impact of model depth, we run GBP Learning with two models: i) a single transposed convolution layer with four filters, and ii) a five layer model comprising stacked transposed convolution and upsampling layers (see App. D). Both the pixel observation and reconstruction factors have robust energies (see Section 5 of Davison:Ortiz:ARXIV2019 for details) to enable the pixel variables to “switch” between being explained by the noisy observation or by the reconstruction model. As a baseline, we compare against a hand-designed method with no learnable parameters. In this pairwise smoothing model, neighbouring pixels are encouraged to have similar intensities via shared smoothness factors (see e.g ortiz2021visual). Robust energies for these factors aid the preservation of edges present in the original image. The baseline also includes robust energies for pixel observation factors, as in our model.
@@
In both the pairwise smoother and GBP Learning, we do inference with GBP. Note that in our models, GBP is jointly estimating the true pixel intensities and the model parameters. We emphasise that our model learns image structure from the noisy images only and without supervision. It is incentivised to do so by the reconstruction factors in noiseless regions. To evaluate our continual learning approach, we use two routines for training the parameters: i) learn them from scratch on each frame and ii) learn them incrementally, by doing filtering on the parameters as described in Section 3.4. The hyperparameters for all models were tuned using the first five frames of the video. Further details of the final models can be found in App. D. Denoising the entire $82$-frame video with the single layer model took $\sim 8$mins on a NVIDIA RTX 3090 GPU, and the five layer model took $\sim 27$mins. Pairwise smoothing took $\sim 2$mins.
@@
The PSNR results are presented in Fig. 3 (and examples of denoised frames in Fig. D1). The models with learnable parameters (orange and green) significantly outperform the classical baseline (blue). We attribute this to the learned models being able to capture image structure, which enables more high-frequency features to be retained while removing the corruption. All methods exhibit a similar relative evolution of PSNR over the course of the video, slowing rising until the fiftieth frame and then falling again. We conjecture that this is due to variations in the content of the images, with some frames being easier to denoise than others.
@@
Figure 3: Video denoising results. Factor graphs with learnable components outperform a hand-specified pairwise smoother. Continual learning of parameters over the video further improves the PSNR over per-frame learning. Shading is ±$\pm 1$ standard error (SE) of $10$ seeds.
@@
We see a clear benefit of depth, with significantly higher PSNR scores for the five layer model (orange/green dotted) over the single layer (orange/green dash). Despite the single layer model having only four filters, we find that it sometimes learns to reconstruct noise when learning continually over the video. To counteract this, we set the prior for each frame to be an interpolation between the previous posterior and the original prior (plotted; see App. D). In contrast, the deeper model does not require additional regularisation in this context. These results imply i) deep factor graphs have better inductive biases, capturing higher-level patterns in the data; ii) GBP Learning is effective in finding aligned multi-layer representations with only local message updates.
@@
In general we see that continual learning leads to an improvement over learning per-frame, suggesting that we can effectively fuse what has been learnt in previous frames and use it to better denoise the current frame.
@@ 5.3 Image Classification
Next, we assess the our method in a supervised learning context, evaluating it on MNIST444yann.lecun.com/exdb/mnist/. MNIST has Creative Commons Attribution-Share Alike 3.0 license. The goal here is to gauge both sample efficiency and the ability to learn continually with only one pass through the training set.
@@
The first baseline allows us to verify our method can do nonlinear learning on images. Further, varying the size of the replay buffer in the second baseline gives an indication of how much information our model can fuse via filtering. All models were tuned on validation sets generated by randomly subsampling $15\%$ of the training set. Further details of the models and hyperparameter selection are included in App. F.
@@
Full dataset training with GBP Learning takes $\sim 3$ hours on NVIDIA RTX3090 GPU. While this is considerably slower than the few minutes to train the CNNs on CPU, we note that the software to train such models has benefited from decades of optimisation, as have well-established GPU hardware platforms. In contrast, our current GBP Learning implementation runs on a hardware and software stack optimised for DL. We believe orders of magnitude efficiency could be gained by running GBP Learning on a more tailored setup including processors with on-chip memory. See Section 6 for discussion.
@@ 5.3.1 Asynchronous Training
To demonstrate the robustness to different update schedules, we also trained in a layer-wise asynchronous manner. At each step, a layer is sampled uniformly, and their messages updated. Such an approach yields a final accuracy of $98.11\pm 0.04\%$ (±$\pm 1$SE over $5$ seeds), close to $98.16\pm 0.03\%$ achieved by the same model trained with synchronous forward/backward sweeps (details and plot of convergence in App. F.4). This result implies that GBP Learning can work well when the model is distributed over multiple processors, without the need for global synchronisation.
@@ 6 Conclusion
xmain.bbl
@@ F.1 Convolutional Factor Graph
Table D4: The convolutional factor graph model used for the MNIST experiment. The dimensions of the layers are chosen such that no padding is necessary. Note that “Recon $\sigma$” denotes the strength of the factors which connect one layer to the next. ${}^{\dagger}$oftmax layer only included at training time, when class observation is available.
@@ F.3 CNN + Replay Buffer Baseline
These hyperparameters were selected based on a validation set of $9,000$ examples, sampled uniformly from the training set. The Adam optimiser kingma2014adam was used for all parameter updates.
@@ F.4 Asynchronous Training
To test the robustness of our method to asynchronous training regimes, we evaluated training the model in Table D4 on MNIST using random layer schedules. At each iteration, we randomly sample $4$ layers, and update the messages within each of the sampled layers in the sampled order. Note that we sample with replacement, meaning that some layers may not be updated at all during that iteration, and some may be updated multiple times.

### v3

@@ 3.3 Efficient GBP
Efficient inference is necessary for our models to be useful in practice. However, the inversion of a $\left(V_{j}-1\right)\times\left(V_{j}-1\right)$ matrix to compute $\Sigma_{\setminus i,\setminus i}^{\left(\phi_{j}+m\right)}$ in (8) has $O\left((V_{j}-1)^{3}\right)$ complexity, bottlenecking GBP. To alleviate this, we exploit the structure of the matrix being inverted. In particular, i) for factors with observation dimension $M=\dim\left(\mathbf{y}\right)<V_{j}$, the precision $\Lambda^{(\phi_{j})}$ (9) is low-rank, and ii) for graphs with scalar variable nodes, $\mathbf{D}$ is diagonal. Thus the sum $\Lambda^{(\phi_{j})}+\mathbf{D}_{\phi_{j}}$ may be efficiently inverted via the Woodbury identity (woodbury1950inverting). Further savings come from reusing intermediates when computing messages to multiple variables (9). These optimisations change the complexity of updating messages from a factor to all $V$ variables, from $O(V\left(V-1\right)^{3})$ to $O\left(VM^{3}\right)$. Space complexity is changed from $O(V^{2})$ to $O(VM+M^{2})$. See App. B for details.
These results constitute significant savings when $M<<V$, raising the question of typical values of observation dimension $M$. For feedforward factors such as (11) and (14), $M$ is equal to the number of output variables connected to the factor. Thus for layers with many outputs, the cubic complexity of the factor to variable update with $M$ may be prohibitive. However, we note that such factors can be decomposed along the output dimension. For example, a dense factor with energy (14) may be decomposed into $M$ smaller factors, one per output variable. As the energy of the original dense factor is recovered by summing the decomposed factor energies, $E_{\mathrm{dense}}=\sum_{j}E_{\mathrm{dense,j}}$, the model is remains unchanged. However, each message update is $M^{2}$ more efficient because one factor with $M$ outputs has been replaced by $M$ factors each with one output.
We now consider how the factor to variable message update complexity translates to the complexity of updating all the messages in a model. As an illustrative example, we consider a model comprising $L$ dense layers, each with $C$ input units and $C$ output units. As described, we can decompose the factor between each pair of layers into $C$ smaller factors, each with $M=1$ and connected to $2\cdot(C+1)$ variables. The complexity of factor to variable message updates in a layer is therefore $O(C^{2})$, the same as the variable to factor message updates. In the non-linear case, we must also account for the computation of the factor Jacobian $\mathbf{J}$ and information vector $\eta$ each time the factor is relinearised. This requires the matrix-vector product $\mathbf{W}_{L}^{\top}\mathrm{vec}\left(\mathbf{X}_{L-1}\right)$, which is also complexity $O(C^{2})$. We thus conclude that the overall complexity for updating all messages in a model of $L$ such layers, with a batch size of $B$, is $O(BLC^{2})$. This is the same complexity as a forward or backwards pass of backpropagation in the equivalent MLP, and the same as that for the approach of lucibello2022deep(lucibello2022deep).
@@ 4 Related Work
Our models can be viewed as probabilistic energy-based models (EBMs; lecun2006tutorial; du2019implicit) whose energy functions are the sum of the quadratic energies for all factors in the graph. The benefit of using a Gaussian factor graph is a model which normalises in closed form without resorting to expensive MCMC sampling. While the quadratic energy may seem constraining, we add capacity with the introduction of non-linear factors and overparameterisation. For most EBMs, parameters are attributes of the factors, where we include them as variables in our graph. As a result, learning and prediction are the same procedure in our approach, but two distinct stages for other EBMs.
Of the EBM family, restricted Boltzmann machines (RBMs smolensky1986information) are of particular relevance. Their factor graph resembles a single layer, fully connected version of our model, however RBMs are models over discrete variables which are unable to capture the statistics of continuous natural images. While exponential family generalisations exist, such as Gaussian-Bernoulli RBMs (welling2004exponential), scaling RBMs to multiple layers remains a challenging problem. To our knowledge there are no working examples of training stacked RBMs jointly, instead they are trained greedily, with each layer learning to reconstruct the activations of the trained and fixed previous layer (hinton2006fast; hinton2006reducing). As an artefact of this, later layers use capacity to learn artificial correlations induced by the early layers, rather than correlations present in the data. Global alignment is then found by fine tuning with either backprop (hinton2006reducing) or wake-sleep (hinton2006fast). We have found GBP Learning to work in multiple layer models without issue.
@@
The factor graphs we consider relate to multi-layer predictive coding (PC) models (millidge2022predictive; rao1999predictive; friston2005theory; buckley2017free; alonso2022theoretical). Designed as a hierarchical model of biological neurons in the brain, multi-layer PC networks are trained by minimising layerwise-local, Gaussian prediction errors similar to our inter-layer consistency factors. Like ours, they also include likelihood factors which ensure observed input/output variables remain close to the observation value. Moreover, some works have noted the suitability of PC for layerwise-parallelism and emerging hardware platforms (salvatori2023brain). Our framework provides a new method to train PC models with GBP, an alternative to the standard approach using gradient descent of a variational free energy (millidge2022predictive). The work of parr2019neuronal(parr2019neuronal) is an exception, and uses BP for inference in a model of biological neurons. However, they assume the model parameters are known, where we infer parameters jointly with activations. Moreover, we consider larger scale machine learning applications. Some PC works perform active inference (buckley2017free; friston2009reinforcement), where an agent may use a PC model to choose actions in uncertain environments. We do not consider this here, but highlight that GBP Learning could be extended for action selection in a similar manner.
Much work has focused on augmenting BP with DL (nachmani2016learning; satorras2021neural; yoon2019inference; lazaro2021query). In contrast, our method does DL within a BP framework. george2017generative(george2017generative) propose a hybrid vision system in which visual features and graph structure are learnt in a separate process to the discrete BP routine used to parse scenes (given features and graph). The work of lazaro2016hierarchical(lazaro2016hierarchical) is more similar to ours in that parameters are treated as variables in the graphical model and updated with BP. However, their factor graph comprises only binary variables, making it ill-suited to natural images. In addition, max-product BP is used to find a MAP solution, and not a marginal posterior estimate. This precludes incremental learning via Bayesian filtering.
@@ 5 Results
We demonstrate our approach with three experiments: some small regression tasks, sequential video denoising and image classification. Our TensorFlow (Tensorflow:MISC2015) implementation is made available222github.com/sethnabarro/gbp_learning/.
@@
Figure 2: GBP Learning in MLP-like factor graphs (2(a)) can solve nonlinear regression and classification tasks. (2(b)) was generated with $8$ hidden units, (2(c)) with $16$ hidden units.
@@ 5.2 Video Denoising
The estimated pixel intensities may explained by the noisy pixel observations and/or the model reconstruction. Both the pixel observation and reconstruction factors have robust energies (see Section 5 of Davison:Ortiz:ARXIV2019(Davison:Ortiz:ARXIV2019)) to enable the pixel variables to “switch” between these explanations.
@@
We evaluate two types of reconstruction model: i) Gaussian factor graphs with learnable parameters, trained with GBP Learning; and ii) a hand-designed baseline, with no learnable parameters. In the learnable models, we examine the impact of model depth by comparing: i) a single transposed convolution layer (factor energies as per (13)) with four filters, and ii) a five layer model comprising transposed convolution and upsampling layers (for details see App. D). In the baseline model, neighbouring pixels are encouraged to have similar intensities via shared smoothness factors (see e.g ortiz2021visual(ortiz2021visual)). We refer to this as “pairwise smoothing”. Robust energies on the smoothness factors aid the preservation of edges present in the original image.
@@
In both the pairwise smoother and GBP Learning, we do inference with GBP. Note that in our models, GBP is jointly estimating the true pixel intensities and the model parameters. We emphasise that our model learns parameters from the noisy images only and without supervision. It is incentivised to do so by the reconstruction factors in noiseless regions. To evaluate our continual learning approach, we try two routines to infer parameters: i) learn them from scratch on each frame and ii) learn them incrementally, by doing filtering on the parameters as described in Section 3.4.
The hyperparameters for all models were tuned using the first five frames of the video. Further details of the final models can be found in App. D. Denoising the entire $82$-frame video with the single layer model took $\sim 8$mins on a NVIDIA RTX 3090 GPU, and the five layer model took $\sim 27$mins. Pairwise smoothing took $\sim 2$mins.
@@
The PSNR results are presented in Fig. 3 (and examples of denoised frames in Fig. D1). All methods exhibit a similar relative evolution of PSNR over the course of the video, slowly rising until the fiftieth frame and then falling again. We conjecture that this is due to varying image content, with some frames being easier to denoise than others.
The models with learnable parameters (orange and green) significantly outperform the classical baseline (blue). We attribute this to the learned models being able to capture image structure, which enables more high-frequency features to be retained while removing the corruption.
@@
Figure 3: Video denoising results. Factor graphs with learnable components outperform a hand-specified pairwise smoother. Continual learning of parameters over the video further improves the PSNR over per-frame learning, and the deep model outperforms the single layer. Shading is ±$\pm 1$ standard error (SE) over $10$ seeds.
@@
We see a clear benefit of depth, with significantly higher PSNR scores for the five layer model (orange/green dotted) over the single layer (orange/green dash). Despite the single layer model having only four filters, we find that it sometimes learns to reconstruct noise when learning continually over the video. To counteract this, we set the prior for each frame to be an interpolation between the previous posterior and the original prior (plotted; see App. D). In contrast, the deeper model does not require additional regularisation when trained continually. These results imply i) deep factor graphs have better inductive biases, capturing higher-level patterns in the data; ii) GBP Learning is can find aligned multi-layer representations with only local message updates.
@@
In general we see that continual learning (green) leads to an improvement over learning per-frame (orange), suggesting that we can effectively fuse what has been learnt in previous frames and use it to better denoise the current frame.
@@ 5.3 Image Classification
Next, we assess our method in a supervised learning context, evaluating it on MNIST444yann.lecun.com/exdb/mnist/. Creative Commons Attribution-Share Alike 3.0 license.. The goal here is to gauge both sample efficiency and the ability to learn continually with only one pass through the training set.
@@
The first baseline allows us to verify our method can do nonlinear learning on images. Further, varying the size of the replay buffer in the second baseline gives an indication of how much information our model can fuse via filtering. All model hyperparameters, including factor strengths $\sigma$, were tuned on validation sets generated by randomly subsampling $15\%$ of the training set. Further details of the models and hyperparameter selection are included in App. F.
@@
Full dataset training with GBP Learning takes $\sim 3$ hours on NVIDIA RTX3090 GPU. While this is much slower than the few minutes to train the CNNs on CPU, we note that the software to train NNs with backprop has benefited from decades of optimisation, as have well-established GPU hardware platforms. In contrast, our current GBP Learning implementation runs on a hardware and software stack optimised for DL. We believe orders of magnitude efficiency could be gained by developing on a more tailored setup with optimised software and hardware with on-chip memory. See Section 6 for discussion.
@@ 5.3.1 Asynchronous Training
To demonstrate the robustness to different update schedules, we also trained in a layer-wise asynchronous manner. At each iteration, we uniformly sample a sequence of $L=4$ layers with replacement to determine the layer update order. Sampling with replacement means some layers may not be updated in an iteration, where some may be updated multiple times. Such an approach yields a final accuracy of $98.11\pm 0.04\%$ (±$\pm 1$SE over $5$ seeds), close to $98.16\pm 0.03\%$ achieved by the same model trained with synchronous forward/backward sweeps (details and plot of convergence in App. F.4). This result implies that GBP Learning can work well when the model is distributed over multiple processors, without the need for global synchronisation.
@@ 5.3.2 Comparison with lucibello2022deep(lucibello2022deep)
We note a significant difference in the relative improvement between FashionMNIST ($0\%$) and CIFAR10 ($11.8\%$), which we conjecture may result from CIFAR10 benefiting more from translation equivariance of the convolutional layer. In FashionMNIST, the objects of interest are mostly centred and fill the frame, where in CIFAR10 the objects are smaller and occur in different regions of the images.
@@ 6 Conclusion
Further, this work has only considered GBP for marginal inference of scalar variables. We note that this is restrictive and prevents capturing the rich correlation structure if the energy landscape of deep networks. By including multidimensional variables in the factor graph, we would expect to mitigate this, and possibly enable more stable BP due to a reduced number of loops in the model. As higher dimensional variables incur greater computational cost per iteration, only variables which are likely to be highly correlated, e.g. the activations or weights within a layer, should be combined.
Acknowledgements
We are grateful to many researchers for helpful discussion relating to this work. In particular, we thank the members of the Dyson Robotics Lab, especially Riku Murai, Joseph Ortiz, Ignacio Alzugaray, Talfan Evans; members of MvdW’s research group, particularly Anish Dhir; and the Vicarious team at Google DeepMind.
SN and AJD are funded by EPSRC Prosperity Partnerships (EP/S036636/1) and Dyson Techonology Ltd.
Impact Statement
Computationally, our method learns via local update rules, which could enable training to be parallelised over arrays of low-power devices in place of GPUs which have high power consumption. In addition, our flexibility to different message schedules could further boost efficiency by executing intra-processor updates at high frequency and inter-processor updates less often, thus reducing communication.
xmain_icml_camera_ready.bbl
@@ F.1 Convolutional Factor Graph
Table D4: The convolutional factor graph model used for the MNIST experiment. The dimensions of the layers are chosen such that no padding is necessary. Note that “Recon $\sigma$” denotes the strength of the factors which connect one layer to the next. Softmax layer only included at training time, when class observation is available. As test time, the final state of the last layer variables after GBP inference is treated as the logit prediction.
@@ F.3 CNN + Replay Buffer Baseline
These hyperparameters were selected based on a validation set of $9,000$ examples, sampled uniformly from the training set. The Adam optimiser (kingma2014adam) was used for all parameter updates.
@@ F.4 Asynchronous Training
To test the robustness of our method to asynchronous training regimes, we evaluated training the model in Table D4 on MNIST using random layer schedules. At each iteration, we uniformly sample $4$ integers in the interval $[1,4]$ with replacement. These integers index the different layers of the network. We then update the messages within each of the sampled layers, in the sampled order. Note that we sample with replacement, meaning that some layers may not be updated at all during a given iteration, and some may be updated multiple times.